What value of "b" will make the system y = 3x +4 and y = 1.5x + b intersect at the point (5,19) ?
Answer by MissionPossible(18)

y = 3x + 4 --------(1)
y = 1.5x + b --------(2)
intersect at the point(5,19) means 5 and 19 should be able to put in the x and y in both eqaution.
But the first equation is useless.
Use only the second equation
Then
19 = (1.5)5 + b
19 = 7.5 + b
Then
b = 19 - 7.5 = 11.5